OpenStudy (anonymous):

please i need help? -find the exact value of the expression given that secx=(3/2) and cscy=3 and x and y are in the quadrant 1: tan(x+y)

OpenStudy (anonymous):

tan(x+y)=(tan(x)+tan(y)) / (1-tan(x)tan(y)) *For tan(x): you got sec(x)=1/cos(x) => you can find cos(x) => use sin^2(x)+cos^2(x)=1, BUT remember to take the positive value of sin(x) since the angle x is in the first quadrant => tan(x)=sin(x)/cos(x) *For tan(y): you got csc(y)=1/sin(y) => you can find sin(y) => use sin^2(y)+cos^2(y)=1 again, BUT remember to take the positive value of cos(y) since the angle y is in the first quadrant => tan(y)=sin(y)/cos(y) Now I think you can do it! :D

OpenStudy (anonymous):

what so i couldnt use the inverse of sec for x and the inverse of csc for y

OpenStudy (anonymous):

*wait

OpenStudy (anonymous):

sorry, i don't really understand what you say... You can go here to see all the trigonometric identities: http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

the values i was given cscy=3 and secx=3/2, why can't i use them directly into the idenity of tan

OpenStudy (anonymous):

you mean tan(x+y)=sin(x+y)/cos(x+y)? Hmm, that's because those angles (x and y) are different from x+y...

OpenStudy (anonymous):

okay so would i have then tan